OpenStudy (anonymous):

Simplify. Assume the divisor is not zero. (8x2y2 – 4xy) ÷ 4xy A. 2xy – 1 B. 2xy C. xy – 1 D. 2xy – y

OpenStudy (anonymous):

\[\frac{ (8 x^{2} y ^{2} -4xy) }{4xy} = \frac{ 8xy }{4xy}\frac{ xy }{ 1 }-\frac{ 4xy }{4xy}\]

OpenStudy (anonymous):

(xy)/(xy) = 1

OpenStudy (anonymous):

So is it B

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@TakeA.I.M. it's A. The first term simplifies to (8/4)xy and the second term is -1. So it's 2xy-1 :)

OpenStudy (anonymous):

oh ok thnx